please help me solve
y=3x-4
-6x+2y=-8

Answer by ankor@dixie-net.com(22740)

please help me solve
y = 3x - 4
Use the above equation for substitution, replace y with (3x-4) in the equation below
-6x + 2y = -8
-6x + 2(3x-4) = -8
-6x + 6x - 8 = -8
Note that we end up
0x = 0
:
There is no unique solution to these two equations (they're the same)